In trangle ABC, the measure of angle B is 15 degree more than three times the measure of angle A. The measure of angle C is 45 degree more than the measure of angle A. Find the mesaure of each angle. ?
Answer by mananth(16946)

let measure of A = x
B will be 3x+15
C= x+45
..
The sum = 180
..
x+3x+15+x+45=180
5x+60=180
5x=180-60
5x=120
x=24 ..... angle A
angle B = 3x+15= 3*24+15
=87 deg
angle C = x+45
=24+45
=69 deg